In an alloy of 100 pounds of zinc and copper there are 75 pounds of copper. How much copper must be added so that the alloy may be 10% zinc ?
Can you please give an equation and detailed explanation

Answer by mananth(16946)

The alloy contains 75 lbs copper in 100 lbs of alloy = 75%
So remaining is Zinc = 25 %
You need to reduce the zinc % to 10% by adding copper which contains no zinc

--- percent ---------------- Amount
Zinc in Alloy 25 ---------------- 100 lbs
copper 0 ---------------- x lbs
Mixture 10 ---------------- 100 + x lbs

25 * 100 + 0 x = 10 ( 100 + x )
2500 + 0 x = 1000 + 10 x
0 x -10 x = -2500 + 1000
-10 x = -1500
/ -10
x = 150 lbs copper

m.ananth@hotmail.ca